a 175 mL sample of carbon dioxide at 678 torr and 295°C is cooled to -3.2°C at constant pressure. what is the final volume of gas in mL?
Answer #1

P = constant

V1 = 175mL

T1 = 295 C = 295+273 =568 K

V2= ???

T2= -3.2 C = -3.2 +273 =269.8K

So using Charles law

V1/T1 = V2/T2

V2 = V1T2/T1 = 175mL x 269.8K / 568K

=83.125 mL